A container is filled to the brim with 1.5 L of mercury at 20°C. As the temperature of the container and mercury is increased to 54°C, a total of 7.8 mL of mercury spill over the brim of the container. Determine the linear expansion coefficient of the material that makes up the container.
Answer #1

Consider the expansion for the mercury :

Vi = initial Volume of mercury = 1.5 L

Vmf = final Volume of mercury = ?

m = linear expansion coefficient of mercury

T = change in temperature = 54 - 20 = 34 oC

final Volume of mercury is given as

Vmf = Vi + 3 VimT    eq-1

Consider the expansion of container :

Vi = initial Volume of container = 1.5 L

Vf = final Volume of container = ?

= linear expansion coefficient of container

T = change in temperature = 54 oC

final Volume of mercury is given as

Vf = Vi + 3 ViT    eq-2

Since 7.8 mL of mercury spill over, hence

Vmf - Vf = 0.0078

( Vi + 3 VimT) - (Vi + 3 ViT) = 0.0078

3 VimT - 3 ViT = 0.0078

3 Vi T (m - ) = 0.0078

(1.5) (34) ((61 x 10-6) - ) = 0.0078

= 10.02 x 10-6